Problem 10-3A

On January 1, 2017, Evers Company purchased the following two machines for use in its production process.
Machine A: The cash price of this machine was $37,500. Related expenditures included: sales tax $2,000, shipping costs $150, insurance during shipping $110, installation and testing costs $120, and $100 of oil and lubricants to be used with the machinery during its first year of operations. Evers estimates that the useful life of the machine is 5 years with a $4,500 salvage value remaining at the end of that time period. Assume that the straight-line method of depreciation is used.
Machine B: The recorded cost of this machine was $180,000. Evers estimates that the useful life of the machine is 4 years with a $10,200 salvage value remaining at the end of that time period.


Which method used to calculate depreciation on Machine B reports the highest amount of depreciation expense in year 1 (2017)? The highest amount in year 4 (2020)? The highest total amount over the 4-year period? Click if you would like to Show Work for this question: Open Show work SHOW LIST OF ACCOUNTS LINK TO TEXT VIDEO: SIMILAR PROBLEM By accessing this Question Assistance, you will learn while you earn points based on the Point Potential Policy set by your instructor
Answer #1

Answer:-

   Before we start solving the given problem as asked in the question let's understand the meaning of "Depreciation"

Whenever we acquired any asset for our business purpose and used that asset in our business to manufacture anything (e.g. Car) , As time passes the value of the Asset start decreasing , It means that asset start losing their monetary value over the period of time. so companies are making provision of certain amount to meet future expenses.

ightarrow While calculating "Depreciation" of any Asset we should have knowledge of the following terms,

1. Cost of Asset

2. Its Useful Life

3. The Scrap Value at the end of its useful life

I'm going to explain the above terms in details,

1. Cost of Asset generally means the actual cost of the asset. It means the amount of money that you pay to take that asset. But The cost of asset also include the following,

Purchase price,

Freight and handling cost,

all directly attributable cost of bringing the asset in working condition'

Delivery cost,

Installation and assembly costs,

Cost Incurred in testing the asset to check the asset is working properly

professional fee pay (if any)

Cost of dismantling or removing the site,

Import duty or any taxes if paid.

ightarrow Useful life generally means the estimated life of the asset.

ightarrowScrap value generally means the value after the estimated useful life of the asset. also said as salvage value.

------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------

Answer of question 1. Machine A

1. Equipment Debit $ 39,880

Cash Credit $ 39,880

Here the cost of lubricant and oil is used for the whole year and the cost of oil and lubricant cannot be depreciated.

2.Depreciation Expense Debit $ 7,076

Accumulated Depreciation Credit $ 7,076   

------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------

Answer of 2nd part of the question. Machine B

1. Use Straight line method of depreciation

The cost of the Asset = $ 39,880

Useful life of the Asset = 5 years

Salvage value or scrap value = $ 4,500

so , to find depreciation under straight line method we use the following formula

depreciation = cost of asset - salvage value div useful life of the asset

= $ 39,880 - $ 4,500 div 5 years = $ 7,076

Depreciation of year 2017 = $ 7,076 , Depreciation of year 2018 = $ 7,076 ,Depreciation of year 2019 = $ 7,076

Depreciation of year 2020 = $ 7076

------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------

Answer of question no 2. machine B

Declining method is well known as written down method i.e. WDV

If we find out depreciation of machine B under SLM than rate will be 25% , under SLM method % of depreciation is 25% and in question it is given that the rate of depreciation is the twice of the SLM (straight line method). so the rate of depreciation is 2*25% = 50%

Note - please find image , working note 1 below of calculation of depreciation.

so, Depreciation in year 2017 = ( $ 180,000 ) * 50% = 90,000

Depreciation of year 2018 = (180,000 - 90,000) * 50% = $ 45,000901000 So 9oRO 82, S00 10,2 00

Depreciation of year 2019 =(180000 - 90,000 - 45,000) * 50% = $ 22,500

Depreciation of year 2020 = ( 180000 - 90,000 - 45,000 - 22500) * 50% = $ 11,250   -------------------------------------------------------------------------------------------------------------------------------------------------------------------------------

Answer of question 3, Machine B

Depreciation under the unit of activity method

Here the useful life of machine in units = 146,000 units and in year 2017 total production = 51,000 unis

so depreciation = (180,000 - 10,200)* 51,000 div 146000 = 59,314

In 2018 = (180,000 - 10,200)* 40500 div 146,000 = 47,102

In 2019 = ( 180,000 - 10,200) * 30,000 div 146,000 = 34890

In 2020 = (180,000 - 10,200) * 24,000 div 146,000 = 27,912

---------------------------------------------------------------------------------------------------------------------------------------------------------------------------

Answer of 3 part of the question.

1. Highest amount of depreciation is under WDV method i.e declining method of $ 90,000

2.Highest amount in year 2020 = $ 27,912

3.Highest total over the 4 year period = $ 90,000 i.e from Declining method.

-----------------------------------------------------------------------------------------------------------------------------------------------------------------------------

We were unable to transcribe this image

We were unable to transcribe this image